OpenStudy (djbreezy):

which is an equation in point-slope form of the line that passes through the points (-4,-1) and (5,7)

OpenStudy (welshfella):

first find the slope of the line which is y2/y1 / x2-x1 (where (x1,y1) and (x2,y2) are the given points.)

OpenStudy (welshfella):

@desiree25 - you have one of the points wrong its (-4,-1) not (-4,1)

OpenStudy (welshfella):

so the slope = (7 - (-1)) / (5-(-4) = 9/9 = 1

OpenStudy (welshfella):

OH * i got it wrong too its 8/9 not 1

OpenStudy (djbreezy):

Sorryyyyyy 1 secc....i have to pit the answer choices

OpenStudy (djbreezy):

A.y+4=9/8(x+1) B. y+1=8/9(x+4) C. y+1=8/7(x+4) D.y-1=8/9(x-4)

OpenStudy (welshfella):

so you rule out A immediately because the slope is 8/9 ot 9/8

OpenStudy (welshfella):

the general equation is y - y1 = m(x - x1) m = 8/9 so we have y - y1 = (8/9) ( x - x1) now all you need do is insert x1 = -4 and y1 = -1 into the above and you'll get the answer

OpenStudy (djbreezy):

i think its D or B

OpenStudy (welshfella):

Remember that 2 negatives together make a plus

OpenStudy (welshfella):

- apply the above rule and you'll get the right one

OpenStudy (welshfella):

- yes - it is B or D

OpenStudy (djbreezy):

I think it B

OpenStudy (welshfella):

-(-1) = +1

OpenStudy (welshfella):

y- (-1) = (8/9) ( x - (-4)) y + 1 = (8/9)( x + 4)

OpenStudy (welshfella):

yes B is correct x - (-1) = x + 1

OpenStudy (djbreezy):

thxxx

OpenStudy (welshfella):

yw brother